# ANLS* - A Universal Document Processing Metric for Generative Large Language Models

### 27 Feb 2024 v2

@@ Abstract
This paper introduces a new metric for generative models called ANLS* for evaluating a wide variety of tasks, including information extraction and classification tasks. The ANLS* metric extends existing ANLS metrics as a drop-in-replacement and is still compatible with previously reported ANLS scores. An evaluation of 7 different datasets and 4 different GLLMs using the ANLS* metric is also provided, demonstrating the importance of the proposed metric.
We also benchmark a novel approach to generate prompts for documents, called SFT, against other prompting techniques such as LATIN. In 15 out of 21 cases, SFT outperforms other techniques and improves the state-of-the-art, sometimes by as much as $15$ percentage points.
@@ 4.2 Experimental Evaluation
In this subsection, the quantitative evaluation of the ANLS* metric on many different datasets and different GLLMs is shown. More precisely, we evaluated two QA datasets (DocVQA [11], MPDocVQA [20]) and five information extraction datasets (Kleister Charity [15], Kleister NDA [15], SROIE [6], VRDU Ad Buy [26], VRDU Registration[26]). The following model versions were used for the evaluation: gpt-3.5-turbo-16k (Version gpt-3.5-turbo-16k-0613), gpt-4-turbo (Version gpt-4-1106-preview) [2], gemini-pro (Version 1.0) [16] and mistral-large (Feb. 2024) [17]. We used LangChain [1] for the implementation with the following system-prompts:
@@
     You are a document information extraction system.  You are given a document and a json with keys that must be extracted from the document.  Here is the document:  {document}  {format_instructions}     Format extractions are automatically generated by LangChain according to the given dataset and keys that must be extracted. For representing the document itself, different prompting methods were evaluated (Simple, LATIN, SFT). More details can be found in the provided source code.
@@
Note that the values for DocLLM are copied from the paper [24] as the model is not available at the time of writing.
Dataset Method gpt-3.5-turbo-16k gpt-4-turbo gemini-pro mistral-large DocLLM*
DocVQA Simple 0.586 0.607 0.586 0.388
Latin Prompt 0.659 0.699 0.676 0.403 0.634
SFT (Ours) 0.809 0.790 0.741 0.540
MPDocVQA Simple 0.348 0.389 0.389 0.239
Latin Prompt 0.413 0.463 0.467 0.289 -
SFT (Ours) 0.547 0.548 0.548 0.377
Kleister Charity Simple 0.490 0.743 0.583 0.534
Latin Prompt 0.442 0.735 0.478 0.540 0.499
SFT (Ours) 0.476 0.763 0.633 0.600
Kleister NDA Simple 0.343 0.695 0.623 0.608
Latin Prompt 0.434 0.705 0.599 0.643 -
SFT (Ours) 0.355 0.703 0.552 0.637
SROIE Simple 0.874 0.835 0.263 0.877
Latin Prompt 0.849 0.851 0.371 0.858 -
SFT (Ours) 0.893 0.873 0.288 0.931
VRDU AdBuy Simple 0.402 0.553 0.510 0.305
Latin Prompt 0.389 0.586 0.556 0.351 -
SFT (Ours) 0.661 0.770 0.685 0.243
VRDU Reg. Simple 0.659 0.676 0.699 0.633
Latin Prompt 0.693 0.673 0.740 0.645 -
SFT (Ours) 0.723 0.711 0.720 0.687
The results are shown in Table 3. It can be seen that special prompt formatting improved scores in almost all cases. Our advanced SFT technique improved scores in 15 out of 21 cases sometimes by a large amount. For example for DocVQA SFT outperformed LATIN by as much as $15$ percentage points. For cases similar to Kleister NDA, where documents contain only text and no positional encoding, we can see that there is no improvement for SFT. Nevertheless, in general documents that contain 2D structure show the importance of specialized prompt formatting techniques. It can also be seen that GPT models are superior to gemini-pro, mistral-large as well as to the custom trained DocLLM model. In one case, gemini-pro is outperformed by 60.5 percentage points. Especially mistral-large struggles with processing documents. We found that the safety filter of gemini-pro is too restrictive and filtered a lot of samples. Finally, it is worth mentioning that gpt-3.5-turbo-16k outperformed the newer gpt-4-turbo model in 2 cases, on DocVQA as well as SROIE.
@@ 5 Conclusion & Discussion
In this paper, we introduce a novel metric called ANLS*, which serves as a plug-in replacement for existing ANLS metrics. This metric is not only applicable for traditional GLLM tasks such as QA, but also for information extraction tasks, and even for more complex outputs. We demonstrate that ANLS* is a versatile metric for a broad range of tasks. We evaluate the ANLS* metric using three different GLLMs on seven different datasets. We also demonstrate that more advanced methods such as SFT outperform prompting techniques, such as LATIN. We found that the GPT models outperform both gemini-pro and the custom-trained DocLLM model. We believe that DocLLM is still too small with 7B parameters to be competitive against GPT models. Unfortunately, it is not available at the time of writing such that combinations of prompting techniques with DocLLM can not be tested yet.

### 21 Mar 2024 v3

@@ Abstract
This paper introduces a new metric for generative models called ANLS* for evaluating a wide variety of tasks, including information extraction and classification tasks. The ANLS* metric extends existing ANLS metrics as a drop-in-replacement and is still compatible with previously reported ANLS scores. An evaluation of 7 different datasets, 6 different GLLMs and 3 different prompting methods using the ANLS* metric is also provided, demonstrating the importance of the proposed metric.
We also benchmark a novel approach to generate prompts for documents, called SFT, against other prompting techniques such as LATIN. In 27 out of 35 cases, SFT outperforms other techniques and improves the state-of-the-art, sometimes by as much as $18$ percentage points.
@@ 4.2 Experimental Evaluation
In this subsection, the quantitative evaluation of the ANLS* metric on many different datasets and different GLLMs is shown. More precisely, we evaluated two QA datasets (DocVQA [11], MPDocVQA [21]) and five information extraction datasets (Kleister Charity [15], Kleister NDA [15], SROIE [6], VRDU Ad Buy [27], VRDU Registration[27]). The following model versions were used for the evaluation: gpt-3.5-turbo-16k (Version gpt-3.5-turbo-16k-0613), gpt-4-turbo (Version gpt-4-1106-preview) [2], gpt-4-vision (Version gpt-4-vision-preview-1106), gemini-pro (Version 1.0) [17], mistral-large (Feb. 2024) [18] as well as claude-3 (Version claude-3-opus) [16].
We used LangChain [1] for the implementation. Whenever a request to a provider failed (e.g. network timeouts etc.), we send the same request with a sleep time of 10 seconds again. After 5 retries we considered the request as failed resulting in ANLS* of 0.0. We used the following prompts to query each model:
@@
     You are a document information extraction system.  You are given a document and a json with keys that must be extracted from the document.  Here is the document:  {document}  {format_instructions}     Format extractions are automatically generated by LangChain according to the given dataset and keys that must be extracted. For representing the document itself, different prompting methods were evaluated (Simple, LATIN, SFT). More details can be found in the provided source code. For vision based models we excluded the document and added one user message per page directly including the corresponding image without executing any OCR.
Note that special prompting techniques were required for mistral-large as well as claude-3 in order to get decent values. We refer to the source code for more details. The results are shown in Table 3.
@@
No document prompting method is required for vision-based models. Therefore Latin Prompting and SFT is not executed for those.
Dataset Method gpt-3.5-turbo-16k gpt-4-turbo gpt-4-vision gemini-pro mistral-large claude-3
DocVQA Simple 0.586 0.607 0.759 0.586 0.445 0.768
Latin Prompting 0.659 0.699 - 0.676 0.447 0.762
SFT (Ours) 0.809 0.790 - 0.741 0.648 0.831
MPDocVQA Simple 0.517 0.635 0.708 0.603 0.364 0.636
Latin Prompting 0.499 0.739 - 0.502 0.335 0.438
SFT (Ours) 0.734 0.781 - 0.616 0.476 0.575
Kleister Charity Simple 0.490 0.743 0.751 0.583 0.652 0.800
Latin Prompting 0.442 0.735 - 0.478 0.576 0.787
SFT (Ours) 0.476 0.763 - 0.633 0.657 0.786
Kleister NDA Simple 0.343 0.695 0.664 0.623 0.637 0.673
Latin Prompting 0.434 0.705 - 0.599 0.624 0.67
SFT (Ours) 0.355 0.703 - 0.552 0.641 0.677
SROIE Simple 0.874 0.835 0.834 0.263 0.855 0.933
Latin Prompting 0.849 0.851 - 0.371 0.863 0.926
SFT (Ours) 0.893 0.873 - 0.288 0.905 0.949
VRDU AD Buy Simple 0.402 0.553 0.640 0.510 0.386 0.577
Latin Prompting 0.389 0.586 - 0.556 0.435 0.608
SFT (Ours) 0.661 0.770 - 0.685 0.594 0.633
VRDU Registration Simple 0.659 0.676 0.665 0.699 0.579 0.685
Latin Prompting 0.693 0.673 - 0.740 0.587 0.715
SFT (Ours) 0.723 0.711 - 0.720 0.639 0.705
GLLM Comparison
First of all, it can be seen that the only model that competes against gpt-4-turbo is claude-3. Both models are good in both, VQA as well as IE tasks. So its not clear yet when to select gpt-4-turbo or claude-3. Nevertheless, it can be seen that gemini-pro, mistral-large as well as gpt-3.5-turbo-16k are quite behind those two models and gpt-4-turbo or claude-3 should be selected for document processing tasks. We believe that Google as well as Mistral will release new competitive versions in the future.
Different Prompting Methods
Table Table 3 shows the result for three different prompting methods. The highest scores of 5 out of 7 datasets were achieved with our SFT method. Only for Kleister Charity and Kleister NDA this was not the case. Results for KleisterNDA are very close ($0.703$ vs. $0.705$). For Kleister Charity we found that documents contain mainly text and almost no positional encoding which may describe why SFT and LATIN did not improve simple prompting. Nevertheless, documents usually contain 2D structures and not purely text which shows the importance of specialized prompt formatting techniques such as SFT.
GLLMs with Vision Capabilities
Prompting techniques are mainly required as those models only support text input. Nevertheless, some models exist that support vision capabilities. Although the gpt-4-vision model reaches decent scores, it is still behind our SFT solution which exploits an OCR scanner. In all cases gpt-4-turbo + SFT outperformed gpt-4-vision. This is a strong indicator that vision models are not yet competitive against specialized OCR scanners when combined with advanced prompting techniques.
Comparison with DocLLM
Wang et al. [25] developed and trained a special generative model for documents with 7B parameters. They report 0.634 on DocVQA and 0.499 zero-shot scores on Kleister Charity. Those scores are quite behind our gpt-4-turbo + SFT solution with 0.79 on DocVQA and 0.763 on Kleister Charity. This shows that smaller specialized models are not yet competitive against large purely text-based models that exploit specialized document prompting techniques.
Open-source Models
We also made tests with mixtral-8x7b, but found that those models are still far behind closed-source models. We, therefore, decided to not include those models in the final evaluation and will report additional results in a future or updated publication.
@@ 5 Conclusion & Discussion
In this paper, we introduce a novel metric called ANLS*, which serves as a plug-in replacement for existing ANLS metrics. This metric is not only applicable for traditional GLLM tasks such as QA, but also for information extraction tasks, and even for more complex outputs. We demonstrate that ANLS* is a versatile metric for a broad range of tasks. We evaluate the ANLS* metric using three different GLLMs on seven different datasets. We also demonstrate that more advanced methods such as SFT outperform prompting techniques, such as LATIN. We found that claude-3 is competitive against gpt-4-turbo, but all others are still behind those two models. For DocLLM we beliefe that this specially trained model is still too small with 7B parameters to be competitive against larger pure text-based models. Unfortunately, it is not available at the time of writing such that combinations of prompting techniques with DocLLM can not be tested yet.